# Dataset Distillation of 3D Point Clouds via Distribution Matching

### v2

@@ 4.1 Experimental Setup
The proposed method was evaluated on the ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets. The ModelNet10, ModelNet40, and ShapeNet are synthetic datasets generated from CAD models, containing 10, 40, and 55 classes, respectively. The ScanObjectNN consists of real-world scanned objects from 15 classes. Following previous methods [29, 28], we measure the classification accuracy trained on the distilled synthetic dataset. We ensure the fairness by training the network 10 times and report the mean and standard deviation of accuracy. We evaluate the performance across different PPC values of 1, 3, and 10, respectively. Each point cloud object contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [14, 15]. The dataset distillation process for the synthetic dataset was optimized for 1,500 iterations, where the performance was evaluated at every 250 iteration by training the PointNet on the synthetic dataset and testing it on the original test dataset.
@@ 4.2.1 Dataset Distillation
We compare the performance of the proposed method with that of 1) the three most representative coreset selection methods: random selection [16], Herding [2, 1], and K-Center [18], 2) two existing image dataset distillation methods of DC [29] and DM [28], and 3) the only existing 3D point clouds dataset distillation method of PCC [27]. Table 1 compares the quantitative performance of the proposed method with that of the existing methods. DM and DC were initialized with random selection. The synthetic datasets were optimized using PointNet [14], and classification performance was evaluated on PointNet to demonstrate their validity. Among the coreset selection methods, Herding achieves the best overall performance. When PPC is set to 1, Herding and K-Center yield identical results since K-Center selects the first data point using the same algorithm to Herding. As PPC increases, Herding consistently outperforms other coreset selection methods, showing its effectiveness. However, since the coreset selection methods rely only on sample selection without further optimization, they struggle to capture complex feature distributions. The features used in DM are extracted from the layer before the classifier, and they discard substantial information through the global max pooling process of PointNet. This information loss weakens the effectiveness of the distribution matching, preventing DM from accurately capturing the feature distribution of the original dataset. DC, on the other hand, aims to match the gradients of the network. However, PointNet is significantly larger than typical networks in image-based distillation tasks, making it difficult for DC to precisely align the gradients. As a result, it consistently underperforms compared to the proposed method. The proposed method outperforms both the coreset selection and existing distillation techniques. When the PPC is small, accurately capturing the original distribution is challenging, leading less noticeable improvement. As the PPC increases, the proposed method more effectively matches the feature distributions significantly improving the performance.
Table 3: Performance comparison across different initialization strategies. The accuracies of three PPC values of 1, 3, and 10 are averaged. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
@@
† DC [29] with Herding initialization corresponds to PCC [27]
Table 4: Comparison of training time (in hours) at the PPC value of 3.
@@
DC 1.52 5.43 7.57 2.19
DM 0.04 0.11 0.15 0.05
Ours 0.08 0.26 0.35 0.11
Table 5: Ablation study of the proposed SADM loss. Experiments were performed at the PPC value of 3. “Random” denotes the results of the random selection [16], and SA is the proposed semantic alignment.
MN10 MN40 SN SONN
@@
4.2.2 Cross-Architecture Generalization
We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [14] and evaluated on different backbone networks including PointNet++[15], DGCNN[22], PointConv [23], and PT [31], respectively. Unlike PointNet, the architectures used for generalization performance comparison follow hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from that of PointNet. This difference makes it essential for the distilled dataset to retain rich and transferable features that generalize well toward different architectures. Due to significant information loss caused by the pooling operation, DM primarily captures the features specific to the PointNet, failing to preserve the original feature distribution. Consequently, the synthetic dataset lacks diverse structural details, resulting in poor generalization performance, particularly on hierarchical models. DC, while performing better than DM, also suffers from key limitations. It matches the gradients obtained from a trained PointNet, the synthetic dataset becomes overfitted to that specific architecture and fails to generalize toward other networks. Consequently, the generalization ability of DC degrades when applied to other architectures. In contrary, the proposed method addresses these limitations by directly matching the feature distributions extracted from randomly initialized network. By comparing feature maps containing richer and less architecture-biased information, the synthetic dataset captures the features relevant to the PointNet while following the original data distribution more closely. Wherease DC suffers from the overfitting and DM loses critical information, the proposed method yields higher adaptability showing strong generalization performance across various backbone networks.
Figure 3: Comparison of synthetic datasets distilled from the ModelNet40 [24] dataset. Point clouds were colorized according to the y-coordinates. From top to bottom, the orginal dataset (first), initialized point cloud models (second), and the distilled synthetic datasets by using DC [29] (third), DM [28] (fourth), and the proposed method (fifth).
4.2.3 Initialization Strategies
Table 5 presents the classification performance on PointNet for DC [29], DM [28], and PCC [27] under different initialization strategies, including uniform noise, random selection [16], Herding [2, 1], and K-Center [18]. Note that the results of PCC are obtained by our implementation, where the PCC is equivalent to performing DC with Herding initialization. When initialized with noise, DC fails to effectively optimize the synthetic dataset because the network has far more parameters than typical networks used for image dataset distillation, making it difficult to align the gradients. Therefore, the original dataset cannot properly guide the training process, leading to poor convergence. In contrast, the structured initializations such as random selection, Herding, or K-Center improve the performance, where the Herding performs the best by selecting the most representative samples. Unlike DC, DM directly matches the feature representations rather than relying on the gradients and hence less sensitive to the initialization strategies. However, due to the information loss, DM consistently underperforms DC even with the structured initializations. In contrast, the proposed method achieves favorable performance, even when initialized from uniform noise. Furthermore, the proposed method significantly outperforms other approaches under the heuristic initialization strategy, such as random selection. This demonstrates that the proposed method is robust with diverse initialization schemes and effectively resolves the instability issues inherent in the existing dataset distillation methods.
4.2.4 Training Time
Table 5 compares the training times between the proposed method and the existing methods. DM [28] achieves the fastest training time as it uses only a small subset of the overall features, however it often ignores many useful features leading to performance degradation. In contrast, DC [29] incurs a significantly higher computational cost as it aligns the gradients across all the parameters. Notably, for the largest dataset of ShapeNet [4], DC requires over 7 hours for training 50 times slower that DM that completes the training in just 9 minutes (0.15 hours). Since the proposed method is based on the distribution matching approach, its training time is marginally increased compared to DM, yet achieves significant improvement in dataset distillation performance.
@@ 4.2.5 Qualitative Comparison of Synthetic Datasets
Figure 3 compares the resulting synthetic datasets distilled by using the proposed method and the existing methods, respectively. The first row shows the original datasets and the second row shows initialized point cloud objects. DC [29] fails to deviate significantly from the initialized objects and causes noise. Similarly, DM [28] also maintains the original shape of the initialized objects while shifting certain points only, that hinders to capture meaningful structural changes. In contrast, as shown in the bottom row, the proposed method successfully preserves the overall semantic structures of 3D objects in each class, while selectively learning essential features. For example, in the airplane class, we observe significant changes in the edges and corners, that are critical for class discrimination, especially around the wings. Also, in the guitar class, the proposed method jointly optimizes the shape and orientation of the synthesized 3D objects.
@@ 4.3.1 Effect of SADM Loss
Table 5 presents the ablation study on the impact of the proposed SADM loss on the classification performance using PointNet [14]. Without preserving the semantic alignment, the unordered nature of point clouds restricts effective feature matching, resulting in nearly identical performance to that of the random selection [16]. That means the model has not been properly optimized. Otherwise, when the semantic alignment is applied, accuracy improves across all datasets, as the aligned featured provide consistent and structured representation that facilitates more effective feature matching during the optimization. This result confirms that preserving the inherent semantic alignment via feature sorting facilitates the network to exploit well-aligned corresponding features and significantly improves the classification performance.
@@ 4.3.2 Effect of Optimal Rotation Estimation
Table 6 presents the ablation study to evaluate the impact of joint optimization of synthetic dataset and rotation angles. We categorized the test datsets into the aligned, mixed, and rotated groups to analyze the effectiveness of the proposed rotation estimation. The aligned group consists of the datasets where the objects maintain consistent orientations, including ModelNet10 and ShapeNet [4]. The mixed group includes ModelNet40 [24] where only certain classes exhibit rotation variations. The rotated group is ScanObjectNN [19] where the objects exhibit arbitrary rotations across all classes. The baseline (first row) shows the performance without optimization which is identical to that of the random selection. In the second row, we present the results where only the orientation of the synthetic data is optimized using the proposed SADM loss, without updating the synthetic dataset. This leads to performance improvements across all the groups compared to random selection. The improvement becomes more significant as the classes with rotation variation become more dominant. Optimizing the synthetic dataset further boosts the performance, with the best results achieved when both the synthetic data and rotation angles are optimized simultaneously. In particular, the accuracy on the rotated dataset improves significantly, demonstrating that the proposed joint optimization enables the model to be resilient to rotation variations.
@@ 4.4 Limitation
While the proposed SADM loss enhances the semantic consistency by sorting the features, it does not completely align semantically meaningful regions across different point cloud objects. Estimating optimal rotations also improves the complexity, particularly when the datasets are already well-aligned to the canonical axes.

### v3

@@ 4.1 Experimental Setup
The proposed method was evaluated on the ModelNet10 [modelnet40], ModelNet40 [modelnet40], ShapeNet [shapenet], and ScanObjectNN [scanobjectnn] datasets. The ModelNet10, ModelNet40, and ShapeNet are synthetic datasets generated from CAD models, containing 10, 40, and 55 classes, respectively. The ScanObjectNN consists of real-world scanned objects from 15 classes. Following previous methods [gm, dm], we measure the classification accuracy trained on the distilled synthetic dataset. We ensure the fairness by training the network 10 times and report the mean and standard deviation of accuracy. We evaluate the performance across different PPC values of 1, 3, and 10. Each point cloud object contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [pointnet, pointnet2]. The dataset distillation process for the synthetic dataset was optimized for 1,500 iterations, where the performance was evaluated at every 250 iterations by training the PointNet on the synthetic dataset and testing it on the original test dataset.
@@ 4.2.1 Dataset Distillation
We compare the performance of the proposed method with that of 1) the three most representative coreset selection methods: random selection [herding(icarl)], Herding [herding(e2e), herding(scail)], and K-Center [kcenter], 2) two existing image dataset distillation methods of DC [gm] and DM [dm]. Table 1 compares the quantitative performance of the proposed method with that of the existing methods. DM and DC were initialized with random selection. The synthetic datasets were optimized using PointNet [pointnet], and classification performance was evaluated on PointNet to demonstrate their validity. Among the coreset selection methods, Herding achieves the best overall performance. When PPC is set to 1, Herding and K-Center yield identical results since K-Center selects the first data point using the same algorithm to Herding. As PPC increases, Herding consistently outperforms other coreset selection methods, showing its effectiveness. However, because coreset selection methods rely solely on sample selection without further optimization, they struggle to capture complex feature distributions. The features used in DM are extracted from the layer before the classifier, and they discard substantial information through the global max pooling process of PointNet. This information loss weakens the effectiveness of the distribution matching, preventing DM from accurately capturing the feature distribution of the original dataset. DC, on the other hand, aims to match the gradients of the network. However, PointNet is significantly larger than typical networks in image-based distillation tasks, making it difficult for DC to precisely align the gradients. As a result, it consistently underperforms compared to the proposed method. The proposed method outperforms both coreset selection and existing distillation techniques. When the PPC is small, accurately capturing the original distribution is challenging, leading less noticeable improvement. As the PPC increases, the proposed method more effectively matches the feature distributions significantly improving the performance.
Table 3: Comparison of cross-dataset generalization performance at the PPC value of 3, where the models are trained on a subset of ModelNet40 and evaluated on a subset of ShapeNet, and vice versa.
Setting DC DM PCC Ours
SN $\rightarrow$ MN40 54.43 49.66 57.43 61.54
MN40 $\rightarrow$ SN 53.75 47.35 49.42 54.40
Table 5: Comparison of training times (in hours) required for dataset distillation at the PPC value of 3.
Method MN10 MN40 SN SONN
DC 1.52 5.43 7.57 2.19
DM 0.04 0.11 0.15 0.05
Ours 0.08 0.26 0.35 0.11
Table 4: Performance comparison across different initialization strategies. The accuracies for PPC values of 1, 3, and 10 are averaged. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
@@
† DC [gm] with Herding initialization corresponds to PCC [pcc]
4.2.2 Cross-Architecture Generalization
We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [pointnet] and evaluated on different backbone networks including PointNet++[pointnet2], DGCNN[dgcnn], PointConv [pointconv], and PT [pointtransformer], respectively. Unlike PointNet, the architectures used for generalization performance comparison follow hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from those of PointNet. This difference makes it essential for the distilled dataset to retain rich and transferable features that generalize well toward different architectures. Due to significant information loss caused by the pooling operation, DM primarily captures the features specific to the PointNet, failing to preserve the original feature distribution. Consequently, the synthetic dataset lacks diverse structural details, resulting in poor generalization performance, particularly on hierarchical models. DC, while performing better than DM, also suffers from key limitations. It matches the gradients obtained from a trained PointNet, the synthetic dataset becomes overfitted to that specific architecture and fails to generalize toward other networks. Consequently, the generalization ability of DC degrades when applied to other architectures. In contrary, the proposed method addresses these limitations by directly matching the feature distributions extracted from randomly initialized network. By comparing feature maps containing richer and less architecture-biased information, the synthetic dataset captures the features relevant to the PointNet while following the original data distribution more closely. Wherease DC suffers from the overfitting and DM loses critical information, the proposed method yields higher adaptability showing strong generalization performance across various backbone networks.
4.2.3 Cross-Dataset Generalization
To further evaluate the generalization capability of the distilled datasets across different domains, we provide an experimental result of cross-dataset experiment in Table 4. Specifically, we constructed the subsets of ShapeNet and ModelNet40, respectively, by selecting the 16 object classes shared between them. Then we performed the experiments in which the dataset was distilled on the subset of ShapeNet and tested on ModelNet40 (SN $\rightarrow$ MN40), and vice versa (MN40 $\rightarrow$ SN). As shown in Table 4, the proposed method consistently achieves the best performance in both settings of the cross-dataset evaluation, indicating that our distilled datasets effectively capture geometric patterns that are useful regardless of the datasets.
4.2.4 Initialization Strategies
Table 4 presents the classification performance of DC [gm], DM [dm], and PCC [pcc] under different initialization strategies, including uniform noise, random selection [herding(icarl)], Herding [herding(e2e), herding(scail)], and K-Center [kcenter]. Note that the results for PCC are obtained using our implementation, where PCC is equivalent to performing DC with Herding initialization. When initialized with noise, DC fails to effectively optimize the synthetic dataset because the network has far more parameters than typical networks used for image dataset distillation, making it difficult to align the gradients. Therefore, the original dataset cannot properly guide the training process, leading to poor convergence. In contrast, the structured initializations such as random selection, Herding, or K-Center improve the performance, where the Herding performs the best by selecting the most representative samples. Unlike DC, DM directly matches feature representations rather than relying on gradients, making it less sensitive to initialization strategies. However, due to information loss, DM consistently underperforms DC, even when structured initializations are used. In contrast, the proposed method achieves favorable performance even when initialized from uniform noise. Furthermore, it significantly outperforms other approaches under heuristic initialization strategies, such as random selection. These results demonstrate that the proposed method is robust across diverse initialization schemes and effectively addresses the instability issues inherent in existing dataset distillation methods.
4.2.5 Training Time
Table 4 compares the training times between the proposed method and the existing methods. DM [dm] achieves the fastest training time as it uses only a small subset of the overall features, however it often ignores many useful features leading to performance degradation. In contrast, DC [gm] incurs a significantly higher computational cost as it aligns the gradients across all the parameters. Notably, for the largest dataset of ShapeNet [shapenet], DC requires over 7 hours for training 50 times slower that DM that completes the training in just 9 minutes (0.15 hours). Since the proposed method is based on the distribution matching approach, its training time is marginally increased compared to DM, yet achieves significant improvement in dataset distillation performance. To further justify the computational overhead, we conducted an experiment where DM was trained with its default setting, whereas the proposed method was trained during the same training time as DM. As shown in Table 7, even under the same training time condition, our method consistently outperforms DM.
Table 6: Performance comparison between DM and the proposed method at the PPC value of 3, where the proposed method was trained during the same time as DM.
@@ 4.2.5 Training Time
DM 77.58 62.04 56.14 23.33
Ours 83.64 70.56 63.34 33.06
Table 7: Ablation study of the proposed semantic alignment (SA) method. Experiments were performed at the PPC value of 3.
Method MN10 MN40 SN SONN
@@ 4.2.6 Qualitative Comparison of Synthetic Datasets
Figure 3 compares the resulting synthetic datasets distilled by using the proposed method and the existing methods, respectively. The first row shows the original datasets and the second row shows initialized point cloud objects. DC [gm] fails to deviate significantly from the initialized objects and causes noise. Similarly, DM [dm] also maintains the original shape of the initialized objects while shifting certain points only, that hinders to capture meaningful structural changes. In contrast, as shown in the bottom row, the proposed method successfully preserves the overall semantic structures of 3D objects in each class, while selectively learning essential features. For example, in the airplane class, we observe significant changes in the edges and corners, that are critical for class discrimination, especially around the wings. Also, in the guitar class, the proposed method jointly optimizes the shape and orientation of the synthesized 3D objects.
Figure 3: Comparison of synthetic datasets distilled from the ModelNet40 [modelnet40] dataset. Point clouds were colorized according to the y-coordinates. From top to bottom, the orginal dataset (first), initialized point cloud models (second), and the distilled synthetic datasets by using DC [gm] (third), DM [dm] (fourth), and the proposed method (fifth).
@@ 4.3.1 Effect of SADM Loss
Table 7 presents an ablation study on the impact of the proposed SADM loss on classification performance using PointNet [pointnet]. Without preserving semantic alignment, the unordered nature of point clouds restricts effective feature matching, resulting in performance nearly identical to that of random selection [herding(icarl)]. This indicates that the model has not been properly optimized. In contrast, when semantic alignment is applied, accuracy improves across all datasets, as the aligned features provide a consistent and structured representation that facilitates more effective feature matching during optimization. These results confirm that preserving inherent semantic alignment via feature sorting enables the network to exploit well-aligned corresponding features, significantly enhancing classification performance.
@@ 4.3.2 Effect of Optimal Rotation Estimation
Table 9 presents the ablation study to evaluate the impact of joint optimization of synthetic dataset and rotation angles. We categorized the test datsets into the aligned, mixed, and rotated groups to analyze the effectiveness of the proposed rotation estimation. The aligned group consists of the datasets where the objects maintain consistent orientations, including ModelNet10 and ShapeNet [shapenet]. The mixed group includes ModelNet40 [modelnet40] where only certain classes exhibit rotation variations. The rotated group is ScanObjectNN [scanobjectnn] where the objects exhibit arbitrary rotations across all classes. The baseline (first row) shows the performance without optimization which is identical to that of the random selection. In the second row, we present the results where only the orientation of the synthetic data is optimized using the proposed SADM loss, without updating the synthetic dataset. This leads to performance improvements across all the groups compared to random selection. The improvement becomes more significant as the classes with rotation variation become more dominant. Optimizing the synthetic dataset further boosts the performance, with the best results achieved when both the synthetic data and rotation angles are optimized simultaneously. In particular, the accuracy on the rotated dataset improves significantly, demonstrating that the proposed joint optimization enables the model to be resilient to rotation variations.
4.3.3 Effect of Point Sorting
We compared the performance of the proposed SADM against two representative point sorting schemes. The axis-aligned sorting method simply orders the points in the ascending order of their $z$-coordinates, without considering any structural or semantic relationships. The Z-order sorting method [morton] assigns an index to each point based on its 3D coordinates simultaneously, ensuring that points close in space receive nearby indices. As shown in Table 9, the existing sorting methods yield only marginal improvements over the baseline that does not use point sorting. Moreover, they fail to provide consistent semantic alignment across different samples. In contrast, the proposed SADM significantly outperforms these methods by achieving semantically consistent matching, thereby enhancing the effectiveness of dataset distillation for 3D point clouds.
@@ 4.3.3 Effect of Point Sorting
Table 9: Performance comparison of different point sorting schemes at the PPC value of 3. For fair comparison, the rotation parameter optimization was not applied.
Method MN10 MN40 SN SONN
Unsorted 75.21 59.96 53.91 20.20
Axis-Aligned 76.44 60.85 53.97 21.80
Z-order [morton] 76.55 61.47 55.17 21.76
SADM 83.17 67.52 60.71 27.66
@@ 4.4 Limitation
While the proposed SADM loss enhances the semantic consistency by sorting the features, it does not completely align semantically meaningful regions across different point cloud objects. Estimating optimal rotations also increases the complexity, particularly when the datasets are already well-aligned to the canonical axes.